HIGH-ACCURACY POTENTIALS FOR VAN DER WAALS SYSTEMS

\underline{RICHARD DAWES}, Missouri University of Science and Technology, Rolla, MO 65409, USA; XIAO-GANG Wang, JAMES BROWN, TUCKER CARRINGTON, JR., Queen’s University, Kingston, Ontario, K7L 3N6 Canada.

Recent experimental studies of vdWs systems including those by Moazzen-Ahmadi and McKellar,1,2 as well as microwave studies by Minei and Novick3,4 have observed previously unknown stable polar isomers for systems such as (NNO)2 and (OCS)2. The multi-welled floppy nature of the PESs and the small barriers between minima place stringent requirements on the PES for a successful theoretical description of these states. An automated method of generating accurate PESs for vdW systems has been developed and is demonstrated here.5,6 A limited number of ab initio data at the explicitly correlated CCSD(T)-F12b level are interpolated into analytic PESs with negligible fitting error. High-accuracy PESs were developed for a number of systems including (NNO)2, (OCS)2, (CO)2, CO2:CS2 and (NH3)2. Using the PESs, the rovibrational Schrdinger equation is solved with a symmetry-adapted Lanczos algorithm and an uncoupled product basis set. All inter-monomer coordinates are included in the calculations. Calculated transition frequencies are in very close agreement with experiment. References (1) M. Dehghani, M. Afshari, Z. Abusara, N. Moazzen-Ahmadi, A. R. W. McKellar, J. Chem. Phys. 126, 164310 (2007). (2) M. Dehghani, M. Afshari, Z. Abusara, N. Moazzen-Ahmadi, A. R. W. McKellar, J. Chem. Phys. 126, 071102 (2007). (3) N. R. Walker, R. Nicholas, A. J. Minei, S. E. Novick, A. C. Legon, J. Mol. Spec. 251, 153 (2008).

(4) A. J. Minei and S. E. Novick, J. Chem. Phys. 126, 101101 (2007). (5) R. Dawes, X.-G. Wang, A. W. Jasper, T. Carrington Jr., J. Chem. Phys. 133, 134304 (2010). (6) X.-G. Wang, T. Carrington Jr., R. Dawes and A. W. Jasper, J. Mol. Spec. 268, 53 (2011).